Question 16 1 pts Consider the market for purple potatoes below and assume that a price ceiling of $30 is imposed by the government. Calculate the deadweight loss: Purple Potato Market $80 $70 $60 $50 $40 $30 $20 $10 S0 500 1000 1500 2000 2500 3000 3500 4000 4500 Pounds of potatoes per day Demand Supply Price Ceiling The deadweight loss is $2500. The deadweight loss is $5000. The deadweight loss is $2000.
Answer #1

Answer
option 2

before price ceiling equilibrium is at Qd=Qs
where
Q=2500 and P=40

The price ceiling create a deadweight loss which is the decrease quantity and price
after price ceiling
Q=Qs=2000
change in quantity =2500-2000=500 units
the willingness to pay at the price ceiling =$50

deadweight loss =0.5* change in quantity *(the willingness to pay at the price ceiling -price ceiling)
=0.5*500*(50-30)
=5000

the deadweight loss is $5000